# ReassembleNet: Learnable Keypoints and Diffusion for 2D Fresco Reconstruction

### v2

@@ 3.4 Semi-Synthetic Dataset creation and Fine-Tuning
To increase training data while reducing the sym-to-real gap of the synthetic data we introduce a large-scale semi-synthetic dataset based on RePAIR [36]. We build from [40], where frescoes are divided into fragments with varied break patterns. The segmentation begins by selecting two points on the fresco’s circumscribed circle. The segmentation line is then randomly divided into straight or curved segments, with curves created using Fourier bases for realistic edges. The generated pieces of this algorithm are not capable of emulating the complexity of fresco datasets. To overcome this limitation, we enhance the approach by incorporating two additional features: applying erosion operation on the piece and applying a slight random rotations and translations. Rather than introducing noise, these adjustments are essential for accurately modeling realistic fragmentation (see Section 5 and Supp. Mat. H for more details).
@@ 4 Evaluation
We evaluate our method on the RePAIR benchmark dataset [36], following the same experimental protocol. The dataset evaluation procedures are detailed in Section 4.1, while the performance of all methods is assessed in Section 4.2. In Section 5, we present an ablation study on the scalability, the selection of the number of keypoints and the semi-synthetic dataset creation. Additional ablation studies on ReassembleNet’s configuration options Supp. Mat. G and results on the semi-synthetic dataset Supp. Mat. H.
@@ 4.1 Dataset, Metrics and Baseline Methods
The dataset serves as a challenging benchmark for testing modern computational and data-driven puzzle-solving methods. It features realistic 2D and 3D fragments of frescos that, due to natural and human-made impacts undergone over time, exhibit erosion, missing pieces, and irregular shapes. The dataset is multi-modal, including high-resolution images and archaeologist-annotated ground truth and metadata. The RePAIR 2D dataset consists of $100$ puzzle samples, with $79$ for training and $21$ for testing. The total number of fragments in RePAIR is $809$, with an average of $8.09$ fragments per puzzle.
Semi-Synthetic Dataset The semi-synthetic dataset has $5000$ samples with a total of $45200$ pieces, resulting in an average of around $9$ pieces per puzzle. Following the RePAIR 2D train-test split ($80$:$20$ ratio), we divided the semi-synthetic dataset into $4000$ training samples and $1000$ testing samples.
Evaluation Metrics. Following [10], we evaluated the methods using the Root Mean Square Error (RMSE) for both translation (in millimeters) and rotation (in degrees), computed with respect to the ground truth. Specifically, we predict the final pose of each keypoint and compute the average polygon translation and rotation as $\boldsymbol{\mu}_{\hat{t}}^{m}=\frac{1}{K}\sum_{i=1}^{K}\hat{\textbf{s}}_{i0}^{m}$ and $\boldsymbol{\mu}_{\hat{r}}^{m}=\frac{1}{K}\sum_{i=1}^{K}\hat{\textbf{r}}_{i0}^{m}$, respectively. We also evaluated the performance of the methods using the $Q_{pos}$ metric [36], which quantifies the overlap between the ground truth fragment poses (translation and rotation) and the reconstructed solution. More details about these metrics can be found in the Supplementary Material E.
@@
Non-learnable approaches: i) the Archaeological Puzzle Solver [4], which applies a greedy ”next best piece” algorithm based on texture. However, they use an outdated extrapolation process, which was replaced by [12] with the stable-diffusion extrapolation method. ii) The Genetic Algorithm reconstruction uses a fitness function based on geometry, specifically the area of the puzzle’s bounding rectangle and the intersection area of overlapping pieces. While perfect solutions minimize both values, this does not guarantee an optimal solution. iii) Greedy geometric matching, which, starting from a random seed fragment, iteratively extends the fragment pose in a greedy fashion based on geometry, in contrast to [4], which relies on texture compatibility.
Learnable approaches: i) DiffAssemble [26] is a GNN-based architecture designed to tackle reassembly tasks using a diffusion model formulation. In this approach, pieces are treated as elements within a set, represented as nodes in a spatial graph. In the 2D scenario, these pieces are modeled as regular patches. ii) PairingNet [40] is a learning-based approach for fragment pair-searching and matching. It uses a graph-based network to extract features, integrates them via a linear transformer module, and employs contrastive loss for global encoding. A weighted fusion module then computes similarity scores to infer adjacent segments.
@@
Non-learnable Derech et al. [4] 0.04 80.96 139.49
Genetic Optimization [36] 0.05 85.63 151.71
Greedy Geom Match [36] 0.02 76.99 135.95
Learnable & No Fine-Tuning DiffAssemble [26] 0.10 131.36 283.55
PairingNet [40] 0.16 98.45 390.43
ReassembleNet w/ no Learnable KP selection 0.34 57.11 16.58
ReassembleNet w/ Frozen Learnable KP selection 0.38 55.72 27.01
ReassembleNet w/ Learnable KP selection 0.22 49.42 21.79
Learnable & Fine-Tuning DiffAssemble [26] 0.10 123.42 280.76
PairingNet [40] 0.13 91.54 364.62
ReassembleNet w/ no Learnable KP selection 0.19 43.51 18.82
ReassembleNet w/ Frozen Learnable KP selection 0.16 40.60 18.39
ReassembleNet w/ Learnable KP selection 0.19 34.28 17.90
Table 1: Results on RePAIR dataset [36] using $Q_{pos}$ from [36] for groundtruth overlap, Root Mean Square Area (RMSE), in terms of, rotation ($\mathcal{R}^{\circ}$) and translation ($\mathcal{T}_{mm}$). Comparing against non-learnable optimization methods [4, 36] and learning (i.e., deep learning) methods [26, 40]. We do not include PuzzleFusion [18] because it runs out-of-memory due to highly irregular shape.
@@ 4.2 RePAIR Dataset Evaluation
Details. We compare ReassembleNet with both non-learnable and learnable methods. We train our model with Adafactor as the optimization algorithm [27] and initialize the learning rate with $0.001$. We set a batch size of $4$. Our method involves selecting $k$ keypoints (sec. 3.2, which is set to $k=20$. For the learnable keypoint selection module, we pre-train to enhance its performance. Further details on the pre-training process are provided in the Supp. Mat. F. We assess the performance of ReassembleNet in the three different configurations for keypoint selection: (i) a strategy based on the no-learnable keypoint selection, (ii) a frozen learnable module, and (iii) a trainable keypoint selection module optimized for the given task.
Results. Table 1 presents the results on the RePAIR dataset. ReassembleNet with learnable keypoint selection and fine-tuning outperforms the other methods. This result emphasizes that representing irregular pieces as points, while selecting the best keypoints as done by ReassembleNet, is effective. It leads to improvements over the second best performing method, Greedy Geom Match [36], by 55% and 86% for RMSE rotation and translation, respectively.
@@
Regarding the results of other methods, DiffAssemble approximates irregular shapes as squared pieces, making it challenging to accurately identify the correct matches. Furthermore, in the original 2D configuration reported in [26], the method relies on a regular grid to arrange the pieces in the output. Since no such grid is present in our case, this introduces additional limitations. PairingNet performs poorly because it employs a pair-matching loss to align contour points, which is unsuitable for puzzle pieces with erosion or gaps between them. Non-learnable approaches, as also stated in [36], achieve low performance, reinforcing the advantages of learned strategies.
@@ 5 Ablation Study
2.5x less memory $10^{0}$$10^{1}$umber of piecesLogarithmic Memory (GB)ReassembleNetNvidia A100 40GB
@@
Table 2 reports the results of the keypoint number evaluation. It is clear that setting $k=20$ yields the best performance while efficiently maximizing the available resources (40 GB of GPU memory).
Semi-Synthetic Dataset Creation Process. As discussed in Section 3.4, we modify the algorithm proposed by [40]. In this experiment, we compare our semi-synthetic dataset creation method with the original algorithm introduced by [40].
@@
[40] 0.18 56.54 26.42
Our 0.19 34.28 17.90
Table 3: Results on the impact of the Semi-Synthetic Dataset Creation process in contrast to [40].
Table 3 reports the results, showing that ReassembleNet trained on our semi-synthetic dataset outperforms ReassembleNet trained on the dataset generated using [40]. This validates the effectiveness of the modifications we introduced to the original algorithm.
@@ Appendix C The Diffusion Process
We define the forward process as a fixed Markov chain that adds noise following a Gaussian distribution to each input, i.e., each keypoints, $\textbf{x}_{i0}^{m}$ to obtain a noisy version, $\textbf{x}_{it}^{m}$, at timestep $t$. Following [16], we adopt the variance $\beta_{t}$ according to a cosine scheduler and define $q(\textbf{x}_{it}^{m}|\textbf{x}_{i0}^{m})$ as:
@@ Appendix D Semi-Synthetic Dataset Evaluation
DiffAssemble [26] 122.92 73.79
PairingNet [40] 60.11 266.84
@@ Results.
Category Method Global Feats Local Feats Geom Feats $Q_{pos}$uparrow RMSE ($\mathcal{R}^{\circ}$) $\downarrow$ RMSE ($\mathcal{T}_{mm}$) $\downarrow$
@@
no Learnable KP selection V V V 0.34 57.11 16.58
Frozen Learnable KP selection V V V 0.38 55.72 27.01
Learnable KP selection V V V 0.22 49.42 21.79
@@
no Learnable KP selection V V V 0.19 43.51 18.82
Frozen Learnable KP selection V V V 0.16 40.60 18.39
Learnable KP selection V V V 0.19 34.28 17.90
@@ Appendix E Metrics Explanation
We also evaluate the performance of the methods using the $Q_{pos}$ metric [36], which quantifies the overlap between the ground truth fragment poses (translation and rotation) and the reconstructed solution. To ensure that the metric is invariant to rigid motions—preventing good solutions from being penalized due to differing global rotations—we first apply a rigid transformation to align the largest reconstructed fragment (referred to as the anchor) with its corresponding ground truth fragment in both translation and rotation. To compute $Q_{pos}$, we first define the area of a fragment, denoted as $A(m)$. In 2D, the shared area can be determined in two different ways: i) by comparing the non-transparent pixels of two large canvases containing all fragments, or ii) by computing the area intersection of the registered 2D point clouds. Additionally, fragments are weighted based on their area, emphasizing the impact of errors on larger fragments. The metric is formally defined as:
@@
where $w_{m}=\frac{\left|A(m)\right|}{\sum_{k=1}^{M}\left|A(k)\right|}$ represents the weight of each fragment, and $A(\tilde{m})$ denotes the area of the fragments with predicted rotation and translation.
@@ Appendix H Qualitative comparison on Semi-Synthetic Dataset Creation Process
In this section, we are reporting a visual representation of the semi-synthetic dataset created following [40] and the final results of the semi-synthetic dataset we were able to create by adding the random erosion of the borders with a slight random rotations and translation.
Figure 6: An illustration of (a) an example of a RePAIR fresco, (b) a synthetic fresco generated using the algorithm proposed by [40], and (c) a synthetic fresco generated using our modified algorithm. The black contour is intentionally added to highlight the borders of the pieces in b and c.
Figure 6 shows the visual differences in the creation of the semi-synthetic dataset. As can be seen, our proposed algorithm (Figure 6(c)) exhibits a certain similarity to Figure 6(a), which is taken from the real-world dataset RePAIR. In contrast, Figure 6(b) clearly shows that the puzzle generated using the algorithm in [40] deviates significantly from the characteristics present in RePAIR: the pieces are assembled to align perfectly without gaps, ensuring a seamless matching between the pieces.
@@ Appendix I More Qualitative Results on RePAIR Dataset
Figure 7: Qualitative results
We report some more qualitative results on the RePAIR dataset. In particular, we report with Figure 7 some failure cases where it can be seen that the model is learning the complexity of groundtruth data.

### v3

@@ 3.4 Semi-Synthetic Dataset Creation and Fine-Tuning
To increase training data while reducing the sym-to-real gap of the synthetic data, we introduce a large-scale semi-synthetic dataset based on RePAIR [tsesmelis2025re]. We build from [zhou2024pairingnet], where frescoes are divided into fragments with varied break patterns. The segmentation begins by selecting two points on the fresco’s circumscribed circle. The segmentation line is then randomly divided into straight or curved segments, with curves created using Fourier bases for realistic edges. The generated pieces of this algorithm are not capable of emulating the complexity of fresco datasets. To overcome this limitation, we enhance the approach by incorporating two additional features: applying erosion operation on the piece and applying a slight random rotations and translations. Rather than introducing noise, these adjustments are essential for accurately modeling realistic fragmentation (see Section 5 and Supp. Mat. H for more details).
@@ 4 Evaluation
We evaluate our method on the RePAIR benchmark dataset [tsesmelis2025re], following the same experimental protocol. The dataset evaluation procedures are detailed in Section 4.1, while the performance of all methods is assessed in Section 4.2. In Section 5, we present an ablation study on the scalability, the selection of the number of keypoints and the semi-synthetic dataset creation. Additional ablation studies on ReassembleNet’s configuration options Supp. Mat. G and results on the semi-synthetic dataset Supp. Mat. H.
@@ 4.1 Dataset, Metrics and Baseline Methods
The dataset serves as a challenging benchmark for testing modern computational and data-driven puzzle-solving methods. It features realistic 2D and 3D fragments of frescos that, due to natural and human-made impacts undergone over time, exhibit erosion, missing pieces, and irregular shapes. The dataset is multi-modal, including high-resolution images and archaeologist-annotated ground truth and metadata. The RePAIR 2D dataset consists of $121$ puzzle samples, with $97$ for training and $24$ for testing. The total number of fragments in RePAIR is $957$, with an average of $7.91$ fragments per puzzle.
Semi-Synthetic Dataset The semi-synthetic dataset has $5000$ samples with a total of $45834$ pieces, resulting in an average of around $9$ pieces per puzzle. Following the RePAIR 2D train-test split (8080:2020 ratio), we divided the semi-synthetic dataset into $4000$ training samples and $1000$ testing samples.
Evaluation Metrics. Following [giuliari2023positional], we evaluated the methods using the Root Mean Square Error (RMSE) for both translation (in millimeters) and rotation (in degrees), computed with respect to the ground truth. Specifically, we predict the final pose of each keypoint and compute the average polygon translation and rotation as $\boldsymbol{\mu}_{\hat{t}}^{m}=\frac{1}{K}\sum_{i=1}^{K}\hat{\textbf{s}}_{i0}^{m}$ and $\boldsymbol{\mu}_{\hat{r}}^{m}=\frac{1}{K}\sum_{i=1}^{K}\hat{\textbf{r}}_{i0}^{m}$, respectively. We also evaluated the performance of the methods using the $Q_{pos}$ metric [tsesmelis2025re], which quantifies the overlap between the ground truth fragment poses (translation and rotation) and the reconstructed solution. More details about these metrics can be found in the Supplementary Material E.
@@
Non-learnable Approaches: i) the Archaeological Puzzle Solver [derech2021solving], which applies a greedy “next best piece” algorithm based on texture. However, they use an outdated extrapolation process, which was replaced by [Harel2024] with the stable-diffusion extrapolation method. ii) The Genetic Algorithm reconstruction uses a fitness function based on geometry, specifically the area of the puzzle’s bounding rectangle and the intersection area of overlapping pieces. While perfect solutions minimize both values, this does not guarantee an optimal solution. iii) Greedy geometric matching, which, starting from a random seed fragment, iteratively extends the fragment pose in a greedy fashion based on geometry, in contrast to [derech2021solving], which relies on texture compatibility.
Learnable Approaches: i) DiffAssemble [scarpellini2024diffassemble] is a GNN-based architecture designed to tackle reassembly tasks using a diffusion model formulation. In this approach, pieces are treated as elements within a set, represented as nodes in a spatial graph. In the 2D scenario, these pieces are modeled as regular patches. ii) PairingNet [zhou2024pairingnet] is a learning-based approach for fragment pair-searching and matching. It uses a graph-based network to extract features, integrates them via a linear transformer module, and employs contrastive loss for global encoding. A weighted fusion module then computes similarity scores to infer adjacent segments.
@@
Non-learnable Derech et al. [derech2021solving] 0.04 80.96 139.49
Genetic Optimization [tsesmelis2025re] 0.05 85.63 151.71
Greedy Geom Match [tsesmelis2025re] 0.02 76.99 135.95
Learnable & No Fine-Tuning DiffAssemble [scarpellini2024diffassemble] 0.10 131.36 283.55
PairingNet [zhou2024pairingnet] 0.16 98.45 390.43
ReassembleNet w/ no Learnable KP selection 0.35 55.01 16.12
ReassembleNet w/ Frozen Learnable KP selection 0.39 51.96 26.67
ReassembleNet w/ Learnable KP selection 0.27 47.61 19.16
Learnable & Fine-Tuning DiffAssemble [scarpellini2024diffassemble] 0.10 123.42 280.76
PairingNet [zhou2024pairingnet] 0.13 91.54 364.62
ReassembleNet w/ no Learnable KP selection 0.16 42.98 18.11
ReassembleNet w/ Frozen Learnable KP selection 0.17 39.12 18.41
ReassembleNet w/ Learnable KP selection 0.21 32.91 17.18
Table 1: Results on RePAIR dataset [tsesmelis2025re] using $Q_{pos}$ from [tsesmelis2025re] for groundtruth overlap, Root Mean Square Error (RMSE), in terms of, rotation ($\mathcal{R}^{\circ}$) and translation ($\mathcal{T}_{mm}$). Comparing against non-learnable optimization methods [derech2021solving, tsesmelis2025re] and learning (i.e., deep learning) methods [scarpellini2024diffassemble, zhou2024pairingnet]. We do not include PuzzleFusion [hosseini2023puzzlefusion] because it runs out-of-memory due to highly irregular shape.
@@ 4.2 RePAIR Dataset Evaluation
Details. We compare ReassembleNet with both non-learnable and learnable methods. We train our model with Adafactor as the optimization algorithm [shazeer2018adafactor] and initialize the learning rate with $0.001$. We set a batch size of $4$. Our method involves selecting $k$ keypoints (sec. 3.2, which is set to $k=20$. For the learnable keypoint selection module, we pre-train to enhance its performance. Further details on the pre-training process are provided in the Supp. Mat. F. We assess the performance of ReassembleNet in the three different configurations for keypoint selection: (i) a strategy based on the no-learnable keypoint selection, (ii) a frozen learnable module, and (iii) a trainable keypoint selection module optimized for the given task.
Results. Table 1 presents the results on the RePAIR dataset. ReassembleNet with learnable keypoint selection and fine-tuning outperforms the other methods. This result emphasizes that representing irregular pieces as points, while selecting the best keypoints as done by ReassembleNet, is effective. It leads to improvements over the second best performing method, Greedy Geom Match [tsesmelis2025re], by 57% and 87% for RMSE rotation and translation, respectively.
@@
Regarding the results of other methods, DiffAssemble approximates irregular shapes as squared pieces, making it challenging to accurately identify the correct matches. Furthermore, in the original 2D configuration reported in [scarpellini2024diffassemble], the method relies on a regular grid to arrange the pieces in the output. Since no such grid is present in our case, this introduces additional limitations. PairingNet performs poorly because it employs a pair-matching loss to align contour points, which is unsuitable for puzzle pieces with erosion or gaps between them. Non-learnable approaches, as also stated in [tsesmelis2025re], achieve low performance, reinforcing the advantages of learned strategies.
PuzzleFusion does not solve the task as its architecture lacks mechanisms to reduce the number of keypoints, therefore, it is computationally intractable on complex geometry pieces.
@@ 5 Ablation Study
[…]
@@
Table 2 reports the results of the keypoint number evaluation. It is clear that setting $k=20$ yields the best performance while efficiently maximizing the available resources (40 GB of GPU memory). The reported VRAM reflects the entire pipeline, including the keypoint selector and two attention modules.
Semi-Synthetic Dataset Creation Process. As discussed in Section 3.4, we modify the algorithm proposed by [zhou2024pairingnet]. In this experiment, we compare our semi-synthetic dataset creation method with the original algorithm introduced by [zhou2024pairingnet].
@@
[zhou2024pairingnet] 0.18 56.54 26.42
Our 0.21 32.91 17.18
Table 3: Results on the impact of the Semi-Synthetic Dataset Creation process in contrast to [zhou2024pairingnet].
Table 3 reports the results, showing that ReassembleNet trained on our semi-synthetic dataset outperforms ReassembleNet trained on the dataset generated using [zhou2024pairingnet]. This validates the effectiveness of the modifications we introduced to the original algorithm.
@@ 6 Conclusion
Acknowledgements
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 964854.
[…]
@@ Appendix C The Diffusion Process
We define the forward process as a fixed Markov chain that adds noise following a Gaussian distribution to each input, i.e., each keypoints, $\textbf{x}_{i0}^{m}$ to obtain a noisy version, $\textbf{x}_{it}^{m}$, at timestep $t$. Following [ho2020denoising], we adopt the variance $\beta_{t}$ according to a cosine scheduler and define $q(\textbf{x}_{it}^{m}|\textbf{x}_{i0}^{m})$ as:
@@
The reverse (denoising) step adds a stochastic term $\sigma_{t}\epsilon_{t}$, where $\epsilon_{t}\sim\mathcal{N}(0,I)$, which governs the randomness injected at each timestep $t$ (see Eq. (11) in [song2020denoising]). By setting $\sigma_{t}=0$, the reverse diffusion becomes fully deterministic.
@@ Appendix D Semi-Synthetic Dataset Evaluation
DiffAssemble [scarpellini2024diffassemble] 122.92 73.79
PairingNet [zhou2024pairingnet] 60.11 266.84
@@ Results.
Category Method Global Feats Local Feats Geom Feats $Q_{pos}$ $\uparrow$ RMSE ($\mathcal{R}^{\circ}$) $\downarrow$ RMSE ($\mathcal{T}_{mm}$) $\downarrow$
@@
no Learnable KP selection V V V 0.35 55.01 16.12
Frozen Learnable KP selection V V V 0.39 51.96 26.67
Learnable KP selection V V V 0.27 47.61 19.16
@@
no Learnable KP selection V V V 0.16 42.98 18.11
Frozen Learnable KP selection V V V 0.17 39.12 18.41
Learnable KP selection V V V 0.21 32.91 17.18
@@ Appendix E Metrics Explanation
We also evaluate the performance of the methods using the $Q_{pos}$ metric [tsesmelis2025re], which quantifies the overlap between the ground truth fragment poses (translation and rotation) and the reconstructed solution. To ensure that the metric is invariant to rigid motions—preventing good solutions from being penalized due to differing global rotations—we first apply a rigid transformation to align the largest reconstructed fragment (referred to as the anchor) with its corresponding ground truth fragment in both translation and rotation. To compute $Q_{pos}$, we first define the area of a fragment, denoted as $A(m)$. In 2D, the shared area can be determined in two different ways: i) by comparing the non-transparent pixels of two large canvases containing all fragments, or ii) by computing the area intersection of the registered 2D point clouds. Additionally, fragments are weighted based on their area, emphasizing the impact of errors on larger fragments. The metric is formally defined as:
@@
where $w_{m}=\frac{\left|A(m)\right|}{\sum_{k=1}^{M}\left|A(k)\right|}$ represents the weight of each fragment, and A(m~)A($\tilde${m}) denotes the area of the fragments with predicted rotation and translation.
Figure 6: An illustration of (a) an example of a RePAIR fresco, (b) a synthetic fresco generated using the algorithm proposed by [zhou2024pairingnet], and (c) a synthetic fresco generated using our modified algorithm. The black contour is intentionally added to highlight the borders of the pieces in b and c.
@@ Appendix H Qualitative comparison on Semi-Synthetic Dataset Creation Process
In this section, we are reporting a visual representation of the semi-synthetic dataset created following [zhou2024pairingnet] and the final results of the semi-synthetic dataset we were able to create by adding the random erosion of the borders with a slight random rotation and translation. Each fragment undergoes morphological erosion using a $3\times 3$ kernel, with between 1-5 iterations randomly simulating varying degrees of degradation. Then, each fragment is randomly augmented with rotation (±3°) or translation (±3 pixels in x and y), applied via affine transformations to introduce geometric variability. These augmentations ensure diversity and realism in the generated dataset.
Figure 6 shows the visual differences in the creation of the semi-synthetic dataset. As can be seen, our proposed algorithm (Figure 6(c)) exhibits a certain similarity to Figure 6(a), which is taken from the real-world dataset RePAIR. In contrast, Figure 6(b) clearly shows that the puzzle generated using the algorithm in [zhou2024pairingnet] deviates significantly from the characteristics present in RePAIR: the pieces are assembled to align perfectly without gaps, ensuring a seamless matching between the pieces.
@@ Appendix I More Qualitative Results on RePAIR Dataset
Figure 7: Qualitative results.
We report some more qualitative results on the RePAIR dataset. In particular, we report with Figure 7 some failure cases where it can be seen that the model is learning the complexity of groundtruth data. We also provide baseline comparison results in Fig. 8.
Figure 8: Qualitative comparison.